[Hallc_running] [New Logentry] RC Update

sjdkay at jlab.org sjdkay at jlab.org
Sun Mar 31 18:20:03 EDT 2019

Logentry Text:
Beam through swing last night was very stable, some small periods of beam down overnight but overall pretty smooth sailing.

We reached (and indeed exceeded) our stats goal for this setting and also acquired the required aluminium dummy target data.

The spectrometers were rotated with assistance from Steve L at 18:00 on Sunday, both spectrometers were moved to their required angles for the next kinematics point, Q2 = 3.0, W = 3.14, x =0.25 low epsilon setting.

We will be at this kinematics setting for quite a long time, the next spectrometer rotation (SHMS moves out 3 degrees to 9.91) will take place roughly at the start of swing shift on April 4th. We'll then be at that setting for a similarly long period of time before moving onto our final Q2 point.

I'll update with how stats are coming tomorrow once the picture becomes clearer.

We will have a meeting tomorrow in the counting house meeting room, details to follow tomorrow. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3674208
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190331/4f9f689c/attachment.html>

More information about the Hallc_running mailing list